OpenStudy (anonymous):

help please A rectangle has sides of length 5 and 8. What is the length of the rectangle's diagonal? Express in radical form.

OpenStudy (anonymous):



OpenStudy (anonymous):

Pythagorean theorem: a^2 + b^2 = c^2 a, b are sides of the triangle c is the hypotenuse The diagonal bisects the rectangle and is therefore the longest side of the resultant triangle. Thus, you put in: 5^2 + 8^2 = c^2 25 + 64 = c^2 c^2 = 89 c = 9.44

OpenStudy (anonymous):

thank you!:)

OpenStudy (anonymous):

Happy to help.

OpenStudy (anonymous):

Please click "Best Response" if you like my description, thanks